# Truncated Affinity Maximization: One-class Homophily Modeling for Graph Anomaly Detection

### v4

@@ 4 Experiments
AUROC iForest 0.3765±${}_{\pm 0.019}$ 0.5118±${}_{\pm 0.018}$ 0.5621±${}_{\pm 0.008}$ 0.5382±${}_{\pm 0.015}$ 0.4363±${}_{\pm 0.020}$ 0.4120±${}_{\pm 0.040}$
ANOMALOUS 0.5652±${}_{\pm 0.025}$ 0.6856±${}_{\pm 0.063}$ 0.4457±${}_{\pm 0.003}$ 0.9021±${}_{\pm 0.005}$ 0.5387±${}_{\pm 0.012}$ 0.4956±${}_{\pm 0.003}$
DOMINANT 0.7590±${}_{\pm 0.010}$ 0.8569±${}_{\pm 0.020}$ 0.5996±${}_{\pm 0.004}$ 0.5677±${}_{\pm 0.002}$ 0.5555±${}_{\pm 0.011}$ 0.4133±${}_{\pm 0.010}$
CoLA 0.7746±${}_{\pm 0.009}$ 0.8233±${}_{\pm 0.001}$ 0.5898±${}_{\pm 0.008}$ 0.8434±${}_{\pm 0.011}$ 0.6028±${}_{\pm 0.007}$ 0.4636±${}_{\pm 0.001}$
SL-GAD 0.8123±${}_{\pm 0.002}$ 0.8479±${}_{\pm 0.005}$ 0.5937±${}_{\pm 0.011}$ 0.7936±${}_{\pm 0.005}$ 0.5677±${}_{\pm 0.005}$ 0.3312±${}_{\pm 0.035}$
HCM-A 0.7980±${}_{\pm 0.004}$ 0.8060±${}_{\pm 0.004}$ 0.3956±${}_{\pm 0.014}$ 0.7387±${}_{\pm 0.032}$ 0.4593 ±${}_{\pm 0.011}$ 0.4593±${}_{\pm 0.005}$
ComGA 0.7683±${}_{\pm 0.004}$ 0.8221±${}_{\pm 0.025}$ 0.5895±${}_{\pm 0.008}$ 0.6055±${}_{\pm 0.000}$ 0.5453±${}_{\pm 0.003}$ 0.4391±${}_{\pm 0.000}$
TAM (Ours) 0.8248±${}_{\pm 0.003}$ 0.8878±${}_{\pm 0.024}$ 0.7064±${}_{\pm 0.010}$ 0.9144±${}_{\pm 0.008}$ 0.6023±${}_{\pm 0.004}$ 0.5643±${}_{\pm 0.007}$
AUPRC iForest 0.0423±${}_{\pm 0.002}$ 0.0372±${}_{\pm 0.001}$ 0.1371±${}_{\pm 0.002}$ 0.0316±${}_{\pm 0.003}$ 0.0269±${}_{\pm 0.001}$ 0.0409±${}_{\pm 0.000}$
ANOMALOUS 0.0652±${}_{\pm 0.005}$ 0.0635±${}_{\pm 0.006}$ 0.0558±${}_{\pm 0.001}$ 0.1898±${}_{\pm 0.004}$ 0.0375±${}_{\pm 0.004}$ 0.0519±${}_{\pm 0.002}$
DOMINANT 0.3102±${}_{\pm 0.011}$ 0.4402±${}_{\pm 0.036}$ 0.1424±${}_{\pm 0.002}$ 0.0314±${}_{\pm 0.041}$ 0.0356±${}_{\pm 0.002}$ 0.0395±${}_{\pm 0.020}$
CoLA 0.3270±${}_{\pm 0.000}$ 0.3235±${}_{\pm 0.017}$ 0.0677±${}_{\pm 0.001}$ 0.2106±${}_{\pm 0.017}$ 0.0449±${}_{\pm 0.002}$ 0.0448±${}_{\pm 0.002}$
SL-GAD 0.3882±${}_{\pm 0.007}$ 0.3784±${}_{\pm 0.011}$ 0.0634±${}_{\pm 0.005}$ 0.1316±${}_{\pm 0.020}$ ${}_{\pm 0.004}$ 0.0350±${}_{\pm 0.000}$
HCM-A 0.3139±${}_{\pm 0.001}$ 0.3413±${}_{\pm 0.004}$ 0.0527±${}_{\pm 0.015}$ 0.0713 ±${}_{\pm 0.004}$ 0.0287±${}_{\pm 0.005}$ 0.0287±${}_{\pm 0.012}$
ComGA 0.3293±${}_{\pm 0.028}$ 0.2873±${}_{\pm 0.012}$ 0.1153±${}_{\pm 0.005}$ 0.0354±${}_{\pm 0.001}$ 0.0374±${}_{\pm 0.001}$ 0.0423±${}_{\pm 0.000}$
TAM (Ours) 0.4182±${}_{\pm 0.005}$ 0.5124±${}_{\pm 0.018}$ 0.2634±${}_{\pm 0.008}$ 0.2233±${}_{\pm 0.016}$ 0.0446±${}_{\pm 0.001}$ 0.0778 ±${}_{\pm 0.009}$
@@ Appendix B Description of Datasets
We conduct the experiments on two real-world dataset with injected anomalies and four real-world with genuine anomalies in diverse online shopping services, social networks, and citation networks, including BlogCatalog [49], ACM [48], Amazon [10], Facebook [56], Reddit, YelpChi [21], as well as four large-scale graph datasets. The statistical information including the number of nodes, edge, the dimension of the feature, and the anomalies rate of the datasets can be found in Tab. 9.
@@
Facebook Social Networks R 4,039 88,234 576 27(0.67%)
Reddit Social Networks R 10,984 175,608 64 366(3.33%)

### v5

@@ 4 Experiments
AUROC iForest 0.3765±0.019 0.5118±0.018 0.5621±0.008 0.5382±0.015 0.4363±0.020 0.4120±0.040
ANOMALOUS 0.5652±0.025 0.6856±0.063 0.4457±0.003 0.9021±0.005 0.5387±0.012 0.4956±0.003
DOMINANT 0.7590±0.010 0.8569±0.020 0.5996±0.004 0.5677±0.002 0.5555±0.011 0.4133±0.010
CoLA 0.7746±0.009 0.8233±0.001 0.5898±0.008 0.8434±0.011 0.6028±0.007 0.4636±0.001
SL-GAD 0.8123±0.002 0.8479±0.005 0.5937±0.011 0.7936±0.005 0.5677±0.005 0.3312±0.035
HCM-A 0.7980±0.004 0.8060±0.004 0.3956±0.014 0.7387±0.032 0.4593 ±0.011 0.4593±0.005
ComGA 0.7683±0.004 0.8221±0.025 0.5895±0.008 0.6055±0.000 0.5453±0.003 0.4391±0.000
TAM (Ours) 0.8248±0.003 0.8878±0.024 0.7064±0.010 0.9144±0.008 0.6023±0.004 0.5643±0.007
AUPRC iForest 0.0423±0.002 0.0372±0.001 0.1371±0.002 0.0316±0.003 0.0269±0.001 0.0409±0.000
ANOMALOUS 0.0652±0.005 0.0635±0.006 0.0558±0.001 0.1898±0.004 0.0375±0.004 0.0519±0.002
DOMINANT 0.3102±0.011 0.4402±0.036 0.1424±0.002 0.0314±0.041 0.0356±0.002 0.0395±0.020
CoLA 0.3270±0.000 0.3235±0.017 0.0677±0.001 0.2106±0.017 0.0449±0.002 0.0448±0.002
SL-GAD 0.3882±0.007 0.3784±0.011 0.0634±0.005 0.1316±0.020 0.0406±0.004 0.0350±0.000
HCM-A 0.3139±0.001 0.3413±0.004 0.0527±0.015 0.0713 ±0.004 0.0287±0.005 0.0287±0.012
ComGA 0.3293±0.028 0.2873±0.012 0.1153±0.005 0.0354±0.001 0.0374±0.001 0.0423±0.000
TAM (Ours) 0.4182±0.005 0.5124±0.018 0.2634±0.008 0.2233±0.016 0.0446±0.001 0.0778 ±0.009
@@ Appendix B Description of Datasets
We conduct the experiments on two real-world dataset with injected anomalies and four real-world with genuine anomalies in diverse online shopping services, social networks, and citation networks, including BlogCatalog [49], ACM [48], Amazon [10], Facebook [56], Reddit, YelpChi [21], as well as four large-scale graph datasets. The statistical information including the number of nodes, edge, the dimension of the feature, and the anomalies rate of the datasets can be found in Tab. LABEL:tab:datasets.
@@
Facebook Social Networks R 1,081 55,104 576 27(2.49%)
Reddit Social Networks R 10,984 168,016 64 366(3.33%)